# Deployability-Centric Infrastructure-as-Code Generation: An LLM-based Iterative Framework

### v1

@@
Abstract
Infrastructure-as-Code (IaC) generation holds significant promise for automating cloud infrastructure provisioning. Recent advances in Large Language Models (LLMs) present a promising opportunity to democratize IaC development by generating deployable infrastructure templates from natural language descriptions, but current evaluation focuses on syntactic correctness while ignoring deployability, the fatal measure of IaC template utility. We address this gap through two contributions: (1) IaCGen, an LLM-based deployability-centric framework that uses iterative feedback mechanism to generate IaC templates, and (2) DPIaC-Eval, a deployability-centric IaC template benchmark consists of 153 real-world scenarios that can evaluate syntax, deployment, user intent, and security. Our evaluation reveals that state-of-the-art LLMs initially performed poorly, with Claude-3.5 and Claude-3.7 achieving only 30.2% and 26.8% deployment success on the first attempt respectively. However, IaCGen transforms this performance dramatically: all evaluated models reach $>$90% passItr@25, with Claude-3.5 and Claude-3.7 achieving 98% success rate. Despite these improvements, critical challenges remain in user intent alignment (25.2% accuracy) and security compliance (8.4% pass rate), highlighting areas requiring continued research. Our work provides the first comprehensive assessment of deployability-centric IaC template generation and establishes a foundation for future research.
Index Terms:
Infrastructure-as-Code, IaC, Code Generation, LLM, AWS CloudFormation
$\dagger$\$$\dagger$\$footnotetext: Equal contribution.**footnotetext: Corresponding author. xiaoyu.sun1@anu.edu.au.
I Introduction
@@
The digital transformation of enterprises has fundamentally altered how organizations manage and deploy computing infrastructure. As cloud adoption accelerates globally, with over 82% of enterprises now utilizing cloud services [1], the complexity and scale of infrastructure management have increased exponentially. Infrastructure-as-Code (IaC) has emerged as the default standard for managing this complexity, enabling organizations to define, provision, and maintain cloud resources through programmable templates that ensure consistency, repeatability, and version control across deployments [2, 3, 4]. This paradigm shift from manual infrastructure management to code-driven automation has become essential for modern DevOps practices [5], supporting the rapid deployment cycles and scalability requirements of contemporary software systems.
@@
However, the widespread adoption of IaC faces significant barriers rooted in its inherent technical complexity and the specialized expertise it demands [6, 7, 8]. Creating effective IaC templates requires a deep understanding of cloud service architectures, resource interdependencies, and platform-specific constraints that extend far beyond traditional programming knowledge [5]. The steep learning curve of IaC languages creates substantial obstacles for development teams lacking dedicated DevOps expertise [9, 10]. This expertise gap not only slows development speed but also increases the risk of misconfigurations that can lead to security vulnerabilities and service outages [11, 12].
@@
Recent advances in Large Language Models (LLMs) present a promising opportunity to democratize IaC development by generating deployable infrastructure templates from natural language descriptions [13, 14, 15]. An effective LLM-based IaC generator could potentially transform infrastructure automation by enabling developers to describe their infrastructure requirements in natural language and receive working deployable IaC in return. This capability would significantly accelerate development workflows and reduce the expertise barrier for cloud adoption.
@@
Despite the promising potential of LLMs in IaC automation, existing research suffers from a critical limitation: an overemphasis on syntactic correctness while neglecting the fundamental requirement of deployability. In operational contexts, deployability represents the primary measure of IaC template quality, as non-deployable templates provide no practical value regardless of syntactic correctness. Furthermore, previous research indicates that LLMs demonstrate considerably lower performance in IaC generation compared to general programming tasks [9, 16], with initial success rates as low as 19%.
@@
To mitigate research gaps, we present DPIaC-Eval, a novel benchmark comprising 153 real-world infrastructure tasks across 58 unique AWS services, with each task including a verified deployable IaC template as ground truth. The formative study shows that mainstream LLMs can only achieve a merely 50% success rate in the simplest scenarios. Therefore, we propose IaCGen, a novel framework that consists of two modules, Iterative IaC Generation and IaC Deployment, integrated through an iterative feedback mechanism that mirrors real DevOps workflows. This design enables models to learn from syntactic and deployment failures, progressively refining their templates.
To evaluate IaCGen, we first assess the effectiveness of IaC template generation with six popular LLMs. The results show that using LLM alone achieves only 20.8$\sim$30.2correctness, while IaCGen achieves 39.0$\sim$66.7%, with 54.6$\sim$91.6correctness within 5 and 10 iterations. Since LLMs exhibit performance differences, we further identify five error patterns and find that different models tend to make different types of error, with the Claude models showing the best performance in avoiding errors. Additionally, we evaluate whether IaCGen can achieve better performance on generating IaC templates with human-in-the-loop feedback. We find that human feedback can improve overall correctness by over 19.5%, help all six LLMs achieve more than 90% passItr@25 accuracy. To verify the trustworthiness of the generated IaC templates, we evaluate their user intent matching with resources and attributes from ground truth, and check the security compliance by Checkov. Over 40% of the templates satisfy resource or attribute requirements, and over 70% meet security compliance by security policies. Furthermore, we discuss the generalizability of IaCGen, its implications for SE-inspired LLM-based frameworks, and the emerging research directions concerning vulnerabilities in IaC templates.
@@
The replication package is available at https://github.com/Tianyi2/IaCGen.
II Background
II-A Infrastructure-as-Code
@@
IaC templates are structured around two fundamental components (resource and parameter). Resources represent the cloud services and infrastructure components to be provisioned. The term “resources” is used instead of “services” because a single cloud service may expose multiple resource types. For instance, Amazon SNS, a message delivery service, provides four distinct resource types (Topic, Subscription, TopicPolicy, and TopicInlinePolicy). In AWS CloudFormation, resources follow the hierarchical naming convention ‘AWS::ServiceName::ResourceType’, clearly delineating the service provider, service name, and specific resource type. Parameters enable template customization by allowing external specification of resource property values. Parameters improve template reusability and flexibility by enabling the same template to be deployed across different environments or use cases with varying configurations. Without parameters, templates would be static and require modification for each deployment scenario. Resource properties define the specific attributes and configurations of infrastructure components, analogous to class attributes in object-oriented programming. For example, an “AWS::SNS::Subscription” resource includes properties such as ‘Endpoint’ (specifying the notification destination) and ‘Protocol’ (defining the delivery method). These properties determine the runtime behavior and integration patterns of the provisioned resources.
Parameters:
  EmailAddress:
    Description: Receiver email address.
    Type: String
    Default: test@example.com
Resources:
  MySubscription:
    Type: ‘AWS::SNS::Subscription’
    Properties:
      Endpoint: !Red EmailAddress
      Protocol: email
The substantial knowledge barrier and widespread adoption of IaC underscore the necessity for automated IaC template generation methods to support developers effectively.
II-B Difference Between IaC Tools
@@
Terraform (TF) and CloudFormation (CF) represent two distinct types of IaC tools, each with unique format language, design philosophy, and operational characteristics. TF uses the HashiCorp Configuration Language (HCL), a declarative language designed with a provider-based architecture that enables multi-cloud deployments across AWS, Azure, Google Cloud, and numerous other platforms [17]. While CF uses JSON or YAML to define resources in a hierarchical structure with deep integration into the AWS’s service ecosystem [18]. Additionally, CF focuses exclusively on AWS, providing native integration with AWS services, security models, and deployment patterns, including features like AWS-specific intrinsic functions and cross-stack references [19].
@@
In addition to the syntactic validity, the program deployability fatally determines real-world IaC template utility. TF’s stateful architecture requires explicit state management and manual resource cleanup procedures [19], making systematic deployability evaluation across hundreds of templates impractical. Also, TF’s lack of native rollback mechanisms means that failed deployments can leave partially created resources in indeterminate states, requiring manual intervention to restore clean testing environments [20, 21, 18]. These operational constraints might explain why existing TF-based benchmark limit evaluation to static analysis [9], leaving the crucial deployability dimension unaddressed.
@@
CF, as AWS’s native IaC language, provides unique methodological advantages that make comprehensive deployability testing feasible. When deploying a CF template, AWS automatically creates a stack to hold the collection of resources defined in the template [22]). Users can create, update, and delete entire resource collections by manipulating stacks, enabling systematic removal of all created resources and restoration of the cloud environment to its initial state through simple stack deletion [22]. Therefore, to explore the automated IaC template generation especially for deployability, we focus on the CloudFormation in this paper.
TABLE I: The criteria of IaC dataset difficulty levels.
Difficulty level LOC Resource Parameter
1 $<$ 50 $<$ 2 $<$ 2
2 $<$ 100 $<$ 4 $<$ 5
3 $<$ 150 $<$ 6 $<$ 9
4 $<$ 200 $<$ 12 $<$ 14
@@ 5 $\geq 200$ $\geq 12$ $\geq 14$
III Benchmark: DPIaC-Eval
Most existing research in code generation has predominantly concentrated on general-purpose programming languages [15, 23, 24, 25], often overlooking the distinct challenges inherent to IaC. Unlike conventional software programs, IaC templates must not only be syntactically correct but also semantically valid and deployable within specific cloud environments. This necessitates a nuanced understanding of cloud services, configurations, interdependencies among resources, and constraints imposed by cloud providers. Thus, to systematically evaluate the effectiveness of IaC template generation, we develop the first deployability-focused benchmark, DPIaC-Eval, comprising 153 diverse CloudFormation IaC templates.
III-A Benchmark Construction
The dataset construction process (Fig. 3) follows a systematic and reproducible workflow designed to collect high-quality AWS CloudFormation templates while ensuring adequacy and ethical compliance. Templates are initially sourced from three primary channels: (1) official AWS documentation and sample libraries, including the AWS CloudFormation Sample Templates repository [26]; (2) repositories within the AWS Samples GitHub organization containing CloudFormation templates [27]; and (3) GitHub repositories identified through a dataset containing URLs of repositories that use CloudFormation templates [28]. Ethical considerations are enforced by confirming appropriate licensing (e.g., MIT, Apache 2.0) before inclusion, ensuring that all templates are legally permissible for research purposes. Following an initial collection of approximately 900 templates, candidates undergo automated pre-processing through a multistage filtering funnel. First, size filtering removes templates that exceed 8,000 tokens or 600 lines, ensuring compatibility with the common LLM generation capacities, resulting in 850 templates. Next, syntax validation verifies templates for structural correctness using cfn-linter [29], yielding 465 valid candidates.
Templates then undergo deployment testing using the AWS SDK boto3 client in a sandbox environment configured with least-privilege IAM roles, reducing the dataset to 200 templates. Finally, a rectification process addresses common deployment failures, including cross-stack dependencies, parameter value completion, missing VPC resources, outdated value references, and dependency ordering adjustments. After rectification, our final dataset comprises 153 high-quality, deployable CloudFormation templates that serve as the foundation for our subsequent IaC generation experiments.
Figure 1: Service and difficulty distribution in Benchmark. The figure only presents the top 16 out of 58 most frequent services and the full list is available in our code repository.
III-B Benchmark Characteristics
Each entry in the benchmark contains two essential elements: (1) a natural language prompt describing the infrastructure requirements and (2) the file path to the ground truth AWS CloudFormation template.
The benchmark comprises 58 commonly used AWS cloud infrastructure services [12, 30, 31]. Fig. 1 illustrates the 16 most frequent services in our dataset, with frequency represented by the number of problems in which each service appears. The distribution spans the full stack of infrastructure components typically required in modern cloud environments: (1) Compute and Serverless: EC2, Lambda, and AutoScaling; (2) Data storage and management: S3, DynamoDB, RDS, and EFS; (3) Networking infrastructure: Virtual Private Cloud (VPC) and API Gateway; (4) Security and Management: Identity and Access Management (IAM), CloudWatch, Logs, and Systems Manager (SSM); (5) Event processing and messaging: EventBridge (Events), Simple Queue Service (SQS), and Simple Notification Service (SNS). This comprehensive coverage ensures our benchmark effectively evaluates LLMs’ capabilities across the diverse range of services essential for real-world cloud infrastructure deployment.
@@ III-B Benchmark Characteristics
The service distribution shown in Fig. 1 reflects real-world usage patterns in production environments, with foundational services like IAM and VPC appearing with higher frequency because they provide the security and networking underpinnings for most cloud architectures. Specialized services demonstrate a proportionally lower representation, mirroring their more targeted usage in production systems.
@@
To facilitate systematic evaluation progression, we classify the benchmark into five difficulty levels. Although acknowledging the inherently subjective nature of drawing boundaries for different difficulty levels, the classification is based on objective metrics including the line of code (LoC) in the IaC template, the number of resources required, and the number of configured parameters. The criteria of difficulty level (1-5) are shown in Table I. As shown in the pie chart in Fig. 1, the benchmark demonstrates a balanced distribution across difficulty levels. This distribution and range ensure the benchmark effectively challenges auomated generation methods in various IaC template scenarios while maintaining real-world relevance.
III-C Comparison with IaC-Eval Benchmark

### v2

# Deployability-Centric Infrastructure-as-Code Generation: Fail, Learn, Refine, and Succeed through LLM-Empowered DevOps Simulation

@@
Abstract.
[…]
Infrastructure-as-Code, IaC, Code Generation, AWS CloudFormation, DevOps, LLM4SE, SE4LLM
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; 2026; Montreal, Canadaisbn: 978-1-4503-XXXX-X/YY/MMccs: Software and its engineering Software creation and management
1. Introduction
@@
The digital transformation of enterprises has fundamentally altered how organizations manage and deploy computing infrastructure. As cloud adoption accelerates globally, with over 82% of enterprises now utilizing cloud services (rackspace-report), the complexity and scale of infrastructure management have increased exponentially. Infrastructure-as-Code (IaC) has emerged as the default standard for managing this complexity, enabling organizations to define, provision, and maintain cloud resources through programmable templates that ensure consistency, repeatability, and version control across deployments (kumara2021s; chinamanagonda2019automating; sokolowski2022infrastructure). This paradigm shift from manual infrastructure management to code-driven automation has become essential for modern DevOps practices (basher2019devops), supporting the rapid deployment cycles and scalability requirements of contemporary software systems.
@@
However, the widespread adoption of IaC faces significant barriers rooted in its inherent technical complexity and the specialized expertise it demands (cusumano2010cloud; qiu2023simplifying; iac_low). Creating IaC templates requires a deep understanding of cloud service architectures, resource interdependencies, and platform-specific constraints that extend far beyond traditional programming knowledge (basher2019devops). The steep learning curve of IaC languages creates substantial obstacles for development teams lacking dedicated DevOps expertise (kon2024iac; hashicorp2023). This expertise gap not only slows the development speed but also increases the risk of misconfigurations that lead to security vulnerabilities and service outages (nasiri2024towards; drosos2024your).
@@
Recent advances in Large Language Models (LLMs) present a promising opportunity to democratize IaC development by generating deployable infrastructure templates from natural language descriptions (chen2021evaluating; liu2023your; liao2024code; liu2025ai). An effective LLM-based IaC generator could potentially transform infrastructure automation by enabling developers to describe their infrastructure requirements in natural language and receive working deployable IaC in return. This capability would significantly accelerate development workflows in DevOps and reduce the expertise barrier for cloud adoption.
@@
However, previous research indicates that LLMs demonstrate considerably lower performance in IaC generation compared to general programming tasks (kon2024iac; ragothaman2024optimizing; han2024chase), with initial success rates as low as 19%. We notice that existing research (kon2024iac; ragothaman2024optimizing; palavalli2024using) suffers from a critical limitation: an overemphasis on syntactic correctness while neglecting the fundamental requirement of deployability. In operational contexts, deployability represents the primary measure of IaC template quality, as non-deployable templates provide no practical value regardless of syntactic correctness.
@@
To mitigate research gaps, we present DPIaC-Eval, a novel benchmark comprising 153 real-world infrastructure tasks across 58 unique AWS services, with expert-authored natural language requirements and verified deployable IaC template as reference templates. The formative study shows that mainstream LLMs can only achieve a merely 50% deployment success rate even in the simplest scenarios (with only one resource and one parameter) in its first attempt. This poor performance stems from fundamental challenges in IaC generation: unlike traditional programming where logic errors are often apparent, IaC’s declarative nature obscures issues until execution, requiring iterative testing and validation. Therefore, we propose IaCGen, a novel deployability-centric IaC generation framework. IaCGen is integrated through an iterative feedback mechanism across three deployability-oriented validation stages, format verification, syntax checking, and live deployment, which mirrors the real DevOps workflows. Such design enables models to learn from syntactic and deployment failures, progressively refining templates.
[…]
We further explore the trustworthiness of the generated IaC templates. First, we evaluate the user intent matching on resources and attributes against the reference template and the 25.2% average coverage indicates critical challenges remain in user intent alignment. Additionally, we utilize Checkov to evaluate the security compliance. The low compliance rate (on applicable policies) of 8.4% indicates a critical need for continued research in this domain. Furthermore, we discuss typical security policy violations (70 distinct security policy violations cross all generated IaC templates), automation potential with human-in-the-loop performance, the implications of DevOps-inspired LLM-based framework, and the threats to validity.
@@
2. Background
2.1. Infrastructure-as-Code
@@
IaC templates are structured around two fundamental components (resource and parameter). Resources represent the cloud services and infrastructure components to be provisioned. The term “resources” is used instead of “services” because a single cloud service may expose multiple resource types. For instance, Amazon SNS, a message delivery service, provides four distinct resource types (Topic, Subscription, TopicPolicy, and TopicInlinePolicy). In AWS CloudFormation, resources follow the hierarchical naming convention “AWS::ServiceName::ResourceType”, clearly delineating the service provider, service name, and specific resource type. Parameters enable template customization by allowing external specification of resource property values. Parameters improve template reusability and flexibility by enabling the same template to be deployed across different environments or use cases with varying configurations. Without parameters, templates would be static and require modification for each deployment scenario. Resource properties define the specific attributes and configurations of infrastructure components, analogous to class attributes in object-oriented programming. For example, as shown in Fig. 1, an “AWS::SNS::Subscription” resource includes properties such as “Endpoint” (specifying the notification destination) and “Protocol” (defining the delivery method). These properties determine the runtime behavior and integration patterns of the provisioned resources. The knowledge barrier and widespread adoption of IaC underscore the need for automated IaC template generation methods to effectively support developers.
2.2. Difference Between IaC Tools
@@
Terraform (TF) and CloudFormation (CF) represent two distinct types of IaC tools, each with unique format language, design philosophy, and operational characteristics. TF uses the HashiCorp Configuration Language (HCL), a declarative language designed with a provider-based architecture that enables multi-cloud deployments across AWS, Azure, Google Cloud, and numerous other platforms (pum2024cloudformation), while CF uses JSON or YAML to define resources in a hierarchical structure with deep integration into the AWS’s service ecosystem (hash-tf-cf-compare). Additionally, CF focuses exclusively on AWS, providing native integration with AWS services, security models, and deployment patterns, including features like AWS-specific intrinsic functions and cross-stack references (geeks-tf-cf-compare).
@@
In addition to the syntactic validity, the program deployability fatally determines real-world IaC template utility. TF’s stateful architecture requires explicit state management and manual resource cleanup procedures (geeks-tf-cf-compare), making systematic deployability evaluation across hundreds of templates impractical. Also, TF’s lack of native rollback mechanisms means that failed deployments can leave partially created resources in indeterminate states, requiring manual cloud resources cleaning (tf-limitation-first; tf-limitation-second; hash-tf-cf-compare). These operational constraints might explain why existing TF-based benchmark limit evaluation to static analysis (kon2024iac), leaving the crucial deployability dimension unaddressed.
@@
CF, as AWS’s native IaC language, provides unique methodological advantages that make comprehensive deployability testing feasible. When deploying a CF template, AWS automatically creates a stack to hold the collection of resources defined in the template (cf-stack). Users can create, update, and delete entire resource collections by manipulating stacks, enabling systematic removal of all created resources and restoration of the cloud environment to its initial state through simple stack deletion (cf-stack). Therefore, to explore the automated IaC template generation especially for deployability, we focus on the CloudFormation in this paper.
(a) Parameters section
(b) Resources section
Figure 1. A sample CloudFormation template creating an SNS email subscription with notification recipient email address defined with parameter.
3. Benchmark: DPIaC-Eval
Most existing research in code generation has predominantly concentrated on general-purpose programming languages (liao2024code; gu2023llm; koziolek2024llm; coignion2024performance), often overlooking the distinct challenges inherent to IaC. Unlike conventional software programs, IaC templates must not only be syntactically correct but also semantically valid and deployable within specific cloud environments. Thus, to systematically evaluate the effectiveness of IaC template generation, we develop the first deployability-focused benchmark, DPIaC-Eval, comprising 153 diverse CloudFormation IaC templates.
3.1. Benchmark Construction
The benchmark construction process (Fig. 2) follows a systematic and reproducible workflow designed to collect high-quality AWS CloudFormation templates and ensure adequacy.
Template Sources. Templates are sourced from two primary channels to ensure diversity and quality: (1) official AWS documentation and sample libraries, including the AWS CloudFormation Sample Templates repository (aws_cfn_templates); (2) GitHub repositories that contain CloudFormation template which we find within the AWS Samples GitHub organization (aws_samples_github) and a dataset containing URLs of repositories that use CloudFormation templates (cirlan2024mining). Ethical considerations are enforced by confirming appropriate licensing (e.g., MIT, Apache 2.0) before inclusion, ensuring that all templates are legally permissible for research purposes.
Preprocessing.  Following an initial collection of approximately 900 templates, candidates undergo automated preprocessing through a multistage filtering funnel. First, size filtering removes templates that exceed 8,000 tokens or 600 lines, ensuring compatibility with the common LLM generation capacities, resulting in 850 templates. Next, syntax validation verifies templates for structural correctness using cfn-linter (cfn-linter), yielding 465 valid candidates. Templates then undergo deployment testing using the AWS SDK (boto3) with a sandbox AWS account configured with least-privilege IAM roles, reducing the dataset to 200 templates. Finally, a rectification process addresses common deployment failures, including cross-stack dependencies, parameter value completion, missing VPC resources, outdated value references, and dependency ordering adjustments. After rectification, our final dataset comprises 153 high-quality, deployable CloudFormation templates that serve as the foundation for our subsequent IaC generation experiments.
Figure 2. Workflow of benchmark DPIaC-Eval construction.
Prompt Construction.  Although CloudFormation templates include a Description section, our analysis of over 900 templates revealed that most omit this section or provide inadequate summaries (e.g., “Deploy Single EC2 Linux Instance” for a 170-line template with complex configurations). This necessitated the manual creation of reasonable natural language descriptions.
For each template, one DevOps practitioner with three years of AWS and CloudFormation experience authored natural language prompts describing the template’s intended functionality. All prompts follow a standardized structure: “We need a CloudFormation template that creates [description of purpose and required services].” The practitioner adhered to specific guidelines to ensure prompt quality: (1) capture essential infrastructure requirements without omitting critical dependencies, (2) reflect realistic practitioner requests that balance specificity with implementation flexibility, and (3) avoid overly prescriptive implementation details.
To enhance prompt quality, we employ DeepSeek-R1 as an editorial assistant, instructed to act as a junior DevOps engineer reviewing the human-authored descriptions alongside original templates. The AI assistant helped refine grammar, identify ambiguities, and suggest missing descriptions. For instance, it identified when prompts omitted critical networking dependencies or contained unclear resource relationships. The original practitioner reviewed all AI suggestions and refined prompts as necessary. For validation, two additional DevOps practitioners with equivalent expertise independently reviewed all prompts. The three practitioners discussed discrepancies and collaboratively finalized the descriptions through consensus. This rigorous validation process achieved an initial inter-rater agreement of 0.68 (Cohen’s Kappa), indicating substantial agreement according to established interpretation guidelines. The resulting prompts remain human-authored and focus on high-level functionality without implementation specifics, mirroring how developers typically specify infrastructure requirements in practice.
(a) AWS service distribution
(b) Difficulty level distribution
Figure 3. Benchmark characteristics. (a) Distribution of AWS services, presenting the top 20 out of 58 most frequent services. The full list is available in the replication package. (b) Distribution of IaC template difficulty levels, showing balanced coverage across five levels.
3.2. Benchmark Characteristics
Each entry in the benchmark contains two essential elements: (1) a natural language prompt describing the infrastructure requirements and (2) the reference CloudFormation template file path. The benchmark comprises 58 commonly used AWS cloud infrastructure services (drosos2024your; jain2023skyplane; stoica2021cloud). Fig. 3(a) illustrates the 20 most frequent services in our dataset, with frequency represented by the number of templates in which each service appears. The distribution spans the full stack of infrastructure components typically required in modern cloud environments: (1) Compute and serverless: EC2, Lambda, and AutoScaling; (2) Data storage and management: S3, DynamoDB, RDS, and Elastic File System (EFS); (3) Networking infrastructure: Virtual Private Cloud (VPC) and API Gateway; (4) Security and management: Identity and Access Management (IAM), CloudWatch, Logs, and Systems Manager (SSM), Key Management Service (KMS), and SecretsManager; (5) Event processing and messaging: EventBridge (Events), Simple Queue Service (SQS), Pipes, and Simple Notification Service (SNS); (6) Data stream processing: Kinesis. This comprehensive coverage ensures our benchmark effectively evaluates LLMs’ capabilities across the diverse range of services essential for real-world cloud infrastructure deployment. The service distribution shown in
Table 1. The criteria of difficulty levels 1-5.
Difficulty Level LoC # Resource # Parameter
1 << 50 << 2 << 2
2 << 100 << 4 << 5
3 << 150 << 6 << 9
4 << 200 << 12 << 14
@@ 5 $\geq 200$ $\geq 12$ $\geq 14$
Fig. 3(a) reflects real-world usage patterns in production environments, with foundational services like IAM and VPC appearing with higher frequency because they provide the security and networking underpinnings for most cloud architectures. Specialized services demonstrate a proportionally lower representation, mirroring their more targeted usage in production systems.
@@
To facilitate systematic evaluation, we classify templates into five difficulty levels, building upon the design established by IaC-Eval (kon2024iac). Acknowledging the inherently subjective nature of drawing boundaries for different difficulty levels, the classification is based on objective metrics including the line of code (LoC) in the IaC template, the number of resources required, and the number of configured parameters. The criteria of difficulty level (1-5) are shown in Table 1. As shown in Fig. 3(b), the benchmark demonstrates a balanced distribution across difficulty levels. This distribution and range ensure the benchmark effectively challenges automated generation methods in various IaC template scenarios while maintaining real-world relevance.
3.3. Comparison with IaC-Eval Benchmark